# Well-posedness of a first-order-in-time model for nonlinear acoustics with nonhomogeneous boundary conditions

### v2

@@ 3.2 Dirichlet-Dirichlet boundary conditions
$f\in L^{2}(0,T;H^{{\sigma-1}})^{d+1},\ \ g\in(H^{{\sigma}})^{d+1},\ h\in H^{1}(0,T;H^{\sigma+\frac{1}{2}}(\partial\Omega))^{d+1}$
@@
    $\mathcal{A}:{D}(-\Delta_{D})^{d+1}\subset H\to H,\ u=(p,v_{1},\dots,v_{d})^{\top}\mapsto(-\zeta\Delta_{D}p,-\mu\Delta_{D}v_{1},\dots,-\mu\Delta_{D}v_{d})^{\top},$
    which is self adjoint, since $\Delta_{D}$ on ${D}(-\Delta_{D})$ is. Since $H^{1}_{0}\hookrightarrow H^{1}$ compactly for $C^{2}$-regular domains and $(L^{2})^{d+1}$ is infinite dimensional, self adjointness is equivalent to the desired spectral decomposition, see [26, Proposition 5.12]. For the smallest eigenvalue $\lambda_{1}$ of $\mathcal{A}$ we have $\lambda_{1}>\min\{\zeta,\mu\}C_{P}^{-1}$. Since ${D}(-\Delta_{D}^{\frac{s}{2}})=(L^{2},H^{2}\cap H^{1}_{0})_{s}$ by Remark 2.1, we obtain ${D}(\mathcal{A}^{\frac{s}{2}})=(\mathbb{H}^{s})^{d+1}$ using the characterization (3.1.1) of $D(-\Delta_{D}^{s})$.
    For $\mathscr{A}:(H^{1}_{0})^{d+1}\subset(L^{2})^{d+1}\to(L^{2})^{d+1},\ u=(p,v)^{\top}\mapsto(\nabla\cdot v,\nabla p)^{\top}$ we have by partial integration, cf. [26, Theorem D.8],
    $\langle\mathscr{A}u,\tilde{u}\rangle=\int_{\Omega}\nabla\cdot v\tilde{p}+\nabla p\cdot\tilde{v}=-\int_{\Omega}\nabla\tilde{p}\cdot v+\nabla\cdot\tilde{v}p+\int_{\partial\Omega}(v\tilde{p}+p\tilde{v})\cdot ndS=\langle u,-\mathscr{A}\tilde{u}\rangle$
    for $u,\tilde{u}\in(H^{1}_{0})^{d+1}={D}(\mathcal{A}^{\frac{1}{2}})\supseteq{D}(\mathcal{A})$. Choosing $u=\tilde{u}$ and using symmetry of $\langle\cdot,\cdot\rangle$ yields the required relation (2.1(ii)).
@@
    Assumption 2.1(iii): We define $\mathfrak{B}$ as in (3). To obtain the estimate (2.1(iii)) note that $H^{s}\hookrightarrow L^{\frac{6}{3-2s}}$ with constant $C_{\hookrightarrow}(s,\Omega)>0$ for $d\leq 3$ and $0\leq s<\frac{3}{2}$, see e.g. [4] and the references therein. So using Hölder’s inequality with $p_{1}=p_{2}=\frac{6}{3-2\sigma},p_{3}=\frac{6}{3-2(1-\sigma)}$ gives
@@
    Assumption 2.2(ii): $\mathfrak{T}_{D}$ is linear and $u\in\tilde{X}$ with $\mathfrak{T}_{D}(u)=0$ implies $u\in X$. This follows, since $(H^{1}_{0})^{d+1}\subset X$ and the fact that functions $u\in H^{1}$ that fulfill $\mathcal{T}(u)=0$ are in $H^{1}_{0}$, c.f [5, Chapter 5.5, Theorem 2].
@@
    Note that estimate (2.1(iii)) also holds for the tilde norms as one can see in (3). Therefore, by the same argument as in the proof of Theorem 2.1 we conclude boundedness of $\tilde{\mathfrak{B}}:\tilde{X}\times\tilde{X}\to Y$.
@@
    Assumption 2.2(iv): To construct the demanded extension $\tilde{h}$ we use the bounded harmonic Dirichlet extension operator $D^{\Delta}:H^{s-\frac{1}{2}}(\partial\Omega)\to H^{s}$ for $s>\frac{1}{2}$ as used in [10]. Defining with $h=(h_{1},...,h_{d+1})^{\top}\in H^{1}(0,T;H^{\sigma+\frac{1}{2}}(\partial\Omega))^{d+1}$
@@
Note that for $\sigma=1$ we obtain the results from [11] if $h=0$. For $\sigma>\frac{1}{2}$ the condition $\tilde{h}(0)-g\in(\mathbb{H}^{{\sigma}})^{d+1}$ translates into $\mathfrak{T}_{D}(h(0))=\mathfrak{T}_{D}(g)$, which is a natural compatibility assumption. In the border case $\sigma=\frac{1}{2}$ we obtain well-posedness with least regular data as possible. In this case there does not exist a trace of the initial condition, but in order to compensate for this, the condition $\tilde{h}(0)-g\in H^{\frac{1}{2}}_{00}$ is required.
@@ 3.3 Neumann-Hodge/Lion boundary conditions
    $\mathcal{A}:{D}(\mathcal{A})\subset H\to H,\ u=(p,v)\mapsto\begin{pmatrix}-\zeta\Delta_{\tilde{N}}p\\ -\mu\Delta_{M}v\end{pmatrix}$
    with ${D}(\mathcal{A})={D}(-\Delta_{N})_{\diagup{C}}\times{D}(-\Delta_{M})$. Since $\Delta_{N}$ and $\Delta_{M}$ are self adjoint, we obtain that $\mathcal{A}$ is also self adjoint. Since $H^{1}_{\diagup{C}}\hookrightarrow L^{2}$ and $H^{1}_{\perp}\hookrightarrow(L^{2})^{d}$ compactly, see [1, Theorem 2.8], we obtain that $\mathcal{A}$ has pure point spectrum, see [26, Proposition 5.12]. Since we exclude constant solutions from $\Delta_{N}u=\lambda u$, we only obtain strictly positive eigenvalues of $\Delta_{N}$. For $\Delta_{M}$ we remark that assuming $\Omega$ is simply connected, yields that there is no eigenvalue $\lambda_{1}$ with $\lambda_{1}=0$, since a Poincaré like inequality holds in ${D}(-\Delta_{M}^{\frac{1}{2}})$, cf [1, Corollary 3.16]. The result that the kernel of $\Delta_{M}$ is trivial if $\Omega$ is simply connected also follows from [20, Theorem 1.2] for less regular domains.
@@
    Assumption 2.1(iii): To show (2.1(iii)) we use the definition of $\mathfrak{B}$ and estimate (3) as obtained in the proof of Theorem 3.1. Then, it remains to show that the $(H^{r})^{d+1}$-norms can be estimated by the $D(\mathcal{A}^{\frac{r}{2}})$ norms. Indeed, we have ${D}(\mathcal{A}^{\frac{s}{2}})\hookrightarrow(H^{s})^{d+1}$, since
    ${D}(\mathcal{A}^{s})=((L^{2})^{d+1},{D}(\mathcal{A}))_{{s}}\hookrightarrow(L^{2},H^{2})_{s}^{d+1}=(H^{2s})^{d+1}$
@@
    The modified embedding (2.1) can be shown as follows. From a Gelfand embedding we have that
    $L^{2}(0,T;(H^{2})^{d})\cap H^{1}(0,T;(L^{2})^{d})\hookrightarrow C(0,T;(H^{1})^{d})$
@@
    $L^{2}(0,T;\mathbb{G}^{2})\cap H^{1}(0,T;(L^{2}_{\diagup{C}})^{*})\hookrightarrow C(0,T;\mathbb{G}^{1}).$
@@
    $L^{2}(0,T;H^{2}_{\diagup{C}})\cap H^{1}(0,T;(L^{2}_{\diagup{C}})^{*})\hookrightarrow C(0,T;H^{1}).$
@@
    Note that estimate (2.1(iii)) also holds for the tilde norms as one can see from (3) and the fact that $H^{2}$ and $H^{2}_{\diagup{C}}$ share the same norm. Therefore, by the same argument as in the proof of Theorem 2.1 we conclude boundedness of $\tilde{\mathfrak{B}}:\tilde{X}\times\tilde{X}\to Y$.
@@
    Assumption 2.2(iv): For the boundary data $h$ we compactly write $h=(h_{1},h_{2},h_{3})$ in the following. To construct the demanded extension $\tilde{h}=(\tilde{h}_{p},\tilde{h}_{v})^{\top}$ we use for $\tilde{h}_{p}$ the Neumann harmonic extension operator $N^{\Delta}:H^{s-\frac{3}{2}}(\partial\Omega)\to H^{s}$, cf. [9], yielding as extension of $h_{1}\in H^{1}(0,T;H^{\frac{1}{2}}(\partial\Omega))$ a unique $\tilde{h}_{p}:=N^{\Delta}h_{1}\in H^{1}(0,T;H^{2}_{\diagup{C}})$. To obtain $\tilde{h}_{v}$, we use Lemma 3.2 with $s=0$, $\chi=0$, $\xi=h_{2}$, $\nu=h_{3}$ giving a unique $\tilde{h}_{v}$ such that
@@ 3.4 Neumann-Dirichlet boundary conditions
$f\in L^{2}(0,T;H^{{\sigma-1}})^{d+1},\ \ g\in(H^{{\sigma}})^{d+1},\ h\in H^{1}(0,T;H^{\sigma-\frac{1}{2}}(\partial\Omega)\times H^{\sigma+\frac{1}{2}}(\partial\Omega)^{d})$
@@
    $\mathcal{A}:{D}(-\Delta_{\tilde{N}})\times{D}(-\Delta_{D})^{d}\subset H\to H,\ u=(p,v_{1},\dots,v_{d})^{\top}\mapsto(-\zeta\Delta_{\tilde{N}}p,-\mu\Delta_{D}v)^{\top},$
@@
Hence, we can apply Theorem 2.2 under the assumptions that $f,g$ are sufficiently small with $f\in L^{2}(0,T;\mathbb{G}^{\sigma-1}\times(\mathbb{H}^{\sigma-1})^{d}),\ g\in\mathbb{G}^{\sigma}\times(\mathbb{H}^{\sigma})^{d}$ to obtain well-posedness of (3.4) if $h=0$. By (3.1.1) and (3.1.2) we obtain the spaces $U=\mathbb{G}^{\sigma+1}\times(\mathbb{H}^{\sigma+1})^{d}$, $W=\mathbb{G}^{\sigma}\times(\mathbb{H}^{\sigma})^{d}$ and $V=\mathbb{G}^{1-\sigma}\times(\mathbb{H}^{1-\sigma})^{d}$, yielding ${X}=L^{2}(0,T;\mathbb{G}^{\sigma+1}\times(\mathbb{H}^{\sigma+1})^{d})\cap H^{1}(0,T;\mathbb{G}^{\sigma-1}\times(\mathbb{H}^{\sigma-1})^{d})$ and $Y=L^{2}(0,T;\mathbb{G}^{\sigma-1}\times(\mathbb{H}^{\sigma-1})^{d}).$
@@
    Assumption 2.2(i): We choose $\tilde{U}=H^{1+\sigma}_{\diagup{C}}\times(H^{1+\sigma})^{d}$, $\tilde{V}=(H^{1-\sigma})^{d+1}$, $\tilde{W}=(H^{\sigma})^{d+1}$ and the required embeddings follow. This gives us $\tilde{X}=L^{2}(0,T;H^{\sigma+1}_{\diagup{C}}\times H^{{\sigma+1}})^{d}\cap H^{1}(0,T;\mathbb{G}^{\sigma-1}\times\mathbb{H}^{{\sigma-1}})^{d}$ and $\tilde{Y}=Y$ from above. The modified embedding (2.1) can again be shown by comining the modified embeddings of the previous proofs of Theorems 3.1, 3.2.
@@ 4 Conclusion
Other types of boundary conditions, such as absorbing boundary conditions, see [27], are also of interest in nonlinear acoustics. It remains an open question whether analogous boundary conditions can be formulated for equation (1).

### v3

@@ 3.2 Dirichlet-Dirichlet boundary conditions
$f\in L^{2}(0,T;H^{{\sigma-1}})^{d+1},\ \ g\in(H^{{\sigma}})^{d+1},\ h\in H^{1}(0,T;H^{\sigma+\frac{1}{2}}(\partial\Omega))^{d+1}$ (49)
@@
    $\mathcal{A}:{D}(-\Delta_{D})^{d+1}\subset H\to H,\ u=(p,v_{1},\dots,v_{d})^{\top}\mapsto(-\zeta\Delta_{D}p,-\mu\Delta_{D}v_{1},\dots,-\mu\Delta_{D}v_{d})^{\top},$ (50)
    which is self adjoint, since $\Delta_{D}$ on ${D}(-\Delta_{D})$ is. Since $H^{1}_{0}\hookrightarrow H^{1}$ compactly for $C^{2}$-regular domains and $(L^{2})^{d+1}$ is infinite dimensional, self adjointness is equivalent to the desired spectral decomposition, see [schmudgen2012, Proposition 5.12]. For the smallest eigenvalue $\lambda_{1}$ of $\mathcal{A}$ we have $\lambda_{1}>\min\{\zeta,\mu\}C_{P}^{-1}$. Since ${D}(-\Delta_{D}^{\frac{s}{2}})=(L^{2},H^{2}\cap H^{1}_{0})_{s}$ by Remark 2.1, we obtain ${D}(\mathcal{A}^{\frac{s}{2}})=(\mathbb{H}^{s})^{d+1}$ using the characterization (43) of $D(-\Delta_{D}^{s})$.
    For $\mathscr{A}:(H^{1}_{0})^{d+1}\subset(L^{2})^{d+1}\to(L^{2})^{d+1},\ u=(p,v)^{\top}\mapsto(\nabla\cdot v,\nabla p)^{\top}$ we have by partial integration, cf. [schmudgen2012, Theorem D.8],
    $\langle\mathscr{A}u,\tilde{u}\rangle=\int_{\Omega}\nabla\cdot v\tilde{p}+\nabla p\cdot\tilde{v}=-\int_{\Omega}\nabla\tilde{p}\cdot v+\nabla\cdot\tilde{v}p+\int_{\partial\Omega}(v\tilde{p}+p\tilde{v})\cdot ndS=\langle u,-\mathscr{A}\tilde{u}\rangle$ (51)
    for $u,\tilde{u}\in(H^{1}_{0})^{d+1}={D}(\mathcal{A}^{\frac{1}{2}})\supseteq{D}(\mathcal{A})$. Choosing $u=\tilde{u}$ and using symmetry of $\langle\cdot,\cdot\rangle$ yields the required relation (10).
@@
    Assumption 2.1(iii): We define $\mathfrak{B}$ as in (41). To obtain the estimate (11) note that $H^{s}\hookrightarrow L^{\frac{6}{3-2s}}$ with constant $C_{\hookrightarrow}(s,\Omega)>0$ for $d\leq 3$ and $0\leq s<\frac{3}{2}$, see e.g. [dinezza2012] and the references therein. So using Hölder’s inequality with $p_{1}=p_{2}=\frac{6}{3-2\sigma},p_{3}=\frac{6}{3-2(1-\sigma)}$ gives
@@
    Assumption 2.2(ii): $\mathfrak{T}_{D}$ is linear and $u\in\tilde{X}$ with $\mathfrak{T}_{D}(u)=0$ implies $u\in X$. This follows, since $(H^{1}_{0})^{d+1}\subset X$ and the fact that functions $u\in H^{1}$ that fulfill $\mathcal{T}(u)=0$ are in $H^{1}_{0}$, c.f [Evans:2010, Chapter 5.5, Theorem 2].
@@
    Note that estimate (11) also holds for the tilde norms as one can see in (3). Therefore, by the same argument as in the proof of Theorem 2.1 we conclude boundedness of $\tilde{\mathfrak{B}}:\tilde{X}\times\tilde{X}\to Y$.
@@
    Assumption 2.2(iv): To construct the demanded extension $\tilde{h}$ we use the bounded harmonic Dirichlet extension operator $D^{\Delta}:H^{s-\frac{1}{2}}(\partial\Omega)\to H^{s}$ for $s>\frac{1}{2}$ as used in [kaltenbacher2011b]. Defining with $h=(h_{1},...,h_{d+1})^{\top}\in H^{1}(0,T;H^{\sigma+\frac{1}{2}}(\partial\Omega))^{d+1}$
@@
Note that for $\sigma=1$ we obtain the results from [QuadraticWave] if $h=0$. For $\sigma>\frac{1}{2}$ the condition $\tilde{h}(0)-g\in(\mathbb{H}^{{\sigma}})^{d+1}$ translates into $\mathfrak{T}_{D}(h(0))=\mathfrak{T}_{D}(g)$, which is a natural compatibility assumption. In the border case $\sigma=\frac{1}{2}$ we obtain well-posedness with least regular data as possible. In this case there does not exist a trace of the initial condition, but in order to compensate for this, the condition $\tilde{h}(0)-g\in H^{\frac{1}{2}}_{00}$ is required.
@@ 3.3 Neumann-Hodge/Lion boundary conditions
    $\mathcal{A}:{D}(\mathcal{A})\subset H\to H,\ u=(p,v)\mapsto\begin{pmatrix}-\zeta\Delta_{\tilde{N}}p\\ -\mu\Delta_{M}v\end{pmatrix}$ (56)
    with ${D}(\mathcal{A})={D}(-\Delta_{N})_{\diagup{C}}\times{D}(-\Delta_{M})$. Since $\Delta_{N}$ and $\Delta_{M}$ are self adjoint, we obtain that $\mathcal{A}$ is also self adjoint. Since $H^{1}_{\diagup{C}}\hookrightarrow L^{2}$ and $H^{1}_{\perp}\hookrightarrow(L^{2})^{d}$ compactly, see [amrouche1998, Theorem 2.8], we obtain that $\mathcal{A}$ has pure point spectrum, see [schmudgen2012, Proposition 5.12]. Since we exclude constant solutions from $\Delta_{N}u=\lambda u$, we only obtain strictly positive eigenvalues of $\Delta_{N}$. For $\Delta_{M}$ we remark that assuming $\Omega$ is simply connected, yields that there is no eigenvalue $\lambda_{1}$ with $\lambda_{1}=0$, since a Poincaré like inequality holds in ${D}(-\Delta_{M}^{\frac{1}{2}})$, cf [amrouche1998, Corollary 3.16]. The result that the kernel of $\Delta_{M}$ is trivial if $\Omega$ is simply connected also follows from [mitrea2004, Theorem 1.2] for less regular domains.
@@
    Assumption 2.1(iii): To show (11) we use the definition of $\mathfrak{B}$ and estimate (3) as obtained in the proof of Theorem 3.1. Then, it remains to show that the $(H^{r})^{d+1}$-norms can be estimated by the $D(\mathcal{A}^{\frac{r}{2}})$ norms. Indeed, we have ${D}(\mathcal{A}^{\frac{s}{2}})\hookrightarrow(H^{s})^{d+1}$, since
    ${D}(\mathcal{A}^{s})=((L^{2})^{d+1},{D}(\mathcal{A}))_{{s}}\hookrightarrow(L^{2},H^{2})_{s}^{d+1}=(H^{2s})^{d+1}$ (57)
@@
    The modified embedding (16) can be shown as follows. From a Gelfand embedding we have that
    $L^{2}(0,T;(H^{2})^{d})\cap H^{1}(0,T;(L^{2})^{d})\hookrightarrow C(0,T;(H^{1})^{d})$ (58)
@@
    $L^{2}(0,T;\mathbb{G}^{2})\cap H^{1}(0,T;(L^{2}_{\diagup{C}})^{*})\hookrightarrow C(0,T;\mathbb{G}^{1}).$ (59)
@@
    $L^{2}(0,T;H^{2}_{\diagup{C}})\cap H^{1}(0,T;(L^{2}_{\diagup{C}})^{*})\hookrightarrow C(0,T;H^{1}).$ (60)
@@
    Note that estimate (11) also holds for the tilde norms as one can see from (3) and the fact that $H^{2}$ and $H^{2}_{\diagup{C}}$ share the same norm. Therefore, by the same argument as in the proof of Theorem 2.1 we conclude boundedness of $\tilde{\mathfrak{B}}:\tilde{X}\times\tilde{X}\to Y$.
@@
    Assumption 2.2(iv): For the boundary data $h$ we compactly write $h=(h_{1},h_{2},h_{3})$ in the following. To construct the demanded extension $\tilde{h}=(\tilde{h}_{p},\tilde{h}_{v})^{\top}$ we use for $\tilde{h}_{p}$ the Neumann harmonic extension operator $N^{\Delta}:H^{s-\frac{3}{2}}(\partial\Omega)\to H^{s}$, cf. [kaltenbacher2011a], yielding as extension of $h_{1}\in H^{1}(0,T;H^{\frac{1}{2}}(\partial\Omega))$ a unique $\tilde{h}_{p}:=N^{\Delta}h_{1}\in H^{1}(0,T;H^{2}_{\diagup{C}})$. To obtain $\tilde{h}_{v}$, we use Lemma 3.2 with $s=0$, $\chi=0$, $\xi=h_{2}$, $\nu=h_{3}$ giving a unique $\tilde{h}_{v}$ such that
@@ 3.4 Neumann-Dirichlet boundary conditions
$f\in L^{2}(0,T;H^{{\sigma-1}})^{d+1},\ \ g\in(H^{{\sigma}})^{d+1},\ h\in H^{1}(0,T;H^{\sigma-\frac{1}{2}}(\partial\Omega)\times H^{\sigma+\frac{1}{2}}(\partial\Omega)^{d})$ (61)
@@
    $\mathcal{A}:{D}(-\Delta_{\tilde{N}})\times{D}(-\Delta_{D})^{d}\subset H\to H,\ u=(p,v_{1},\dots,v_{d})^{\top}\mapsto(-\zeta\Delta_{\tilde{N}}p,-\mu\Delta_{D}v)^{\top},$ (62)
@@
Hence, we can apply Theorem 2.2 under the assumptions that $f,g$ are sufficiently small with $f\in L^{2}(0,T;\mathbb{G}^{\sigma-1}\times(\mathbb{H}^{\sigma-1})^{d}),\ g\in\mathbb{G}^{\sigma}\times(\mathbb{H}^{\sigma})^{d}$ to obtain well-posedness of (3.4) if $h=0$. By (43) and (45) we obtain the spaces $U=\mathbb{G}^{\sigma+1}\times(\mathbb{H}^{\sigma+1})^{d}$, $W=\mathbb{G}^{\sigma}\times(\mathbb{H}^{\sigma})^{d}$ and $V=\mathbb{G}^{1-\sigma}\times(\mathbb{H}^{1-\sigma})^{d}$, yielding ${X}=L^{2}(0,T;\mathbb{G}^{\sigma+1}\times(\mathbb{H}^{\sigma+1})^{d})\cap H^{1}(0,T;\mathbb{G}^{\sigma-1}\times(\mathbb{H}^{\sigma-1})^{d})$ and $Y=L^{2}(0,T;\mathbb{G}^{\sigma-1}\times(\mathbb{H}^{\sigma-1})^{d}).$
@@
    Assumption 2.2(i): We choose $\tilde{U}=H^{1+\sigma}_{\diagup{C}}\times(H^{1+\sigma})^{d}$, $\tilde{V}=(H^{1-\sigma})^{d+1}$, $\tilde{W}=(H^{\sigma})^{d+1}$ and the required embeddings follow. This gives us $\tilde{X}=L^{2}(0,T;H^{\sigma+1}_{\diagup{C}}\times H^{{\sigma+1}})^{d}\cap H^{1}(0,T;\mathbb{G}^{\sigma-1}\times\mathbb{H}^{{\sigma-1}})^{d}$ and $\tilde{Y}=Y$ from above. The modified embedding (16) can again be shown by comining the modified embeddings of the previous proofs of Theorems 3.1, 3.2.
@@ 4 Conclusion
All of the well-posedness results are established under the assumption that the data are sufficiently small. If the solution possesses enough regularity, the embedding in 16 is not required, and global well-posedness can be inferred, as shown in [QuadraticWave] for homogeneous Dirichlet boundary conditions. It is expected that these results carry over to the case of inhomogeneous boundary conditions as well.
@@
Other types of boundary conditions, such as absorbing boundary conditions, see [Shevchenko2015], are also of interest in nonlinear acoustics. It remains an open question whether analogous boundary conditions can be formulated for equation (1).
Concerning the limit of vanishing damping coefficients $\zeta$ and $\mu$ in (1), note that, as seen in the energy estimate (2.1.2), the nonlinear terms must be controlled using the regularity provided by the symmetric operator $\mathcal{A}$. Consequently, for low regularity data, the controllability of the nonlinear terms relies on the Laplace operator. In contrast, if the initial data are sufficiently regular, well-posedness for the case $\zeta=\mu=0$ can be established using hyperbolic methods. This scenario, among other aspects, is studied in the unfinished work [LehnerMeliani2026].